How can I find secant theta if tangent theta equals 8/15 and theta terminates in quadrant three?
Answer by lwsshak3(11628)

How can I find secant theta if tangent theta equals 8/15 and theta terminates in quadrant three?
**
O=opposite side
A=adjacent side
H=hypotenuse
..
tan=O/A=8/15 (in quadrant III where sec<0
O=8
H=15
A=√(H^2-O^2)=√(225-64)=√161
sec=H/A
=15/√161
=(15√161)/161